OpenStudy (anonymous):

Solve 3x 2 + 2 = 11 - x 2

OpenStudy (anonymous):

You mean: \[3x^2 + 2 = 11 - x^2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Firstly add \(x^2\) both the sides first.. and tell me what did you get?

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

Just add \(x^2\) will lead you to: \[3x^2 + x^2 + 2 = 11 - x^2 + x^2\] Can you solve here for Right Hand Side??

OpenStudy (anonymous):

That would be 11

OpenStudy (anonymous):

Yep: And \(3x^2 + x^2 = ??\)

OpenStudy (anonymous):

That's 0

OpenStudy (anonymous):

I dont know I did it on the calculator

OpenStudy (anonymous):

\(3x^2 + x^2\) = ??

OpenStudy (anonymous):

Let \(x^2 = Banana\).. See, if you have 3 bananas and I give you one more bananas then how many you have now???

OpenStudy (anonymous):

thats 4

OpenStudy (anonymous):

Yep, so \(4 \; bananas = 4x^2\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So, now we have: \[4x^2 + 2 = 11\] Now subtract 2 from both the sides: \[4x^2 + 2 - 2 = 11 - 2\] Can you solve it further??

OpenStudy (anonymous):

What will be left on LHS??

OpenStudy (anonymous):

16x^2

OpenStudy (anonymous):

2 -2 = ??

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Yep: and 11- 2 = ??

OpenStudy (anonymous):

9

OpenStudy (anonymous):

So you are left with: \[4x^2 + \cancel{2} - \cancel{2} = 11 - 2 \implies 4x^2 = 9\] Now divide by 4 here: \[\frac{4x^2}{4} = \frac{9}{4}\]

OpenStudy (anonymous):

In LHS, what will get cancelled??

OpenStudy (anonymous):

4x^2

OpenStudy (anonymous):

and 4 on the right

OpenStudy (anonymous):

Only 4 will cancel out there: \[\frac{\cancel{4}x^2}{\cancel{4}} = \frac{9}{4}\] So you are left with: \[x^2 = \frac{9}{4}\] Take square root now and you will get: \[\sqrt{x^2} = \sqrt{\frac{9}{4}}\] And: \[x = \sqrt{\frac{9}{4}}\] Solve the RHS to find the answer..

OpenStudy (anonymous):

What is \(\sqrt{9}\) = ??

OpenStudy (anonymous):

And after that what is \(\sqrt{4}\) = ??

OpenStudy (anonymous):

3 Thank You very much 2

OpenStudy (anonymous):

SO: \[\color{green}{x = \frac{3}{2}}\]

OpenStudy (anonymous):

You are welcome dear..

OpenStudy (anonymous):

You can add \(\pm\) in front of 3/2 if you want..